### v5

# Denoising Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

Abstract
We propose denoising diffusion variational inference (DDVI), a black-box variational inference algorithm for latent variable models which relies on diffusion models as flexible approximate posteriors. Specifically, our method introduces an expressive class of diffusion-based variational posteriors that perform iterative refinement in latent space; we train these posteriors with a novel regularized evidence lower bound (ELBO) on the marginal likelihood inspired by the wake-sleep algorithm. Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks. We find that DDVI improves inference and learning in deep latent variable models across common benchmarks as well as on a motivating task in biology—inferring latent ancestry from human genomes—where it outperforms strong baselines on 1000 Genomes dataset.
1 Introduction
We are interested in amortized black-box variational inference problems of the form
$\log\;p_{\boldsymbol{\theta}}(\mathbf{x})\geq$ $\max_{\boldsymbol{\phi}}\;\;\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\;p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})-\log\;q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})\right]$
$:=\max_{\boldsymbol{\phi}}\;\;\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}}),$ (1)
in which we approximate the marginal likelihood $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ of a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ with an evidence lower bound $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ that is a function of an approximate variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$. We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks, topic models (blei2003latent), variational autoencoders (VAEs), and broad classes of $p_{\boldsymbol{\theta}}$ defined via modern probabilistic programming frameworks (gordon2014probabilistic).
Maximizing $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ over $\boldsymbol{\phi}$} yields a variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that approximates $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ as well as a tight bound on $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ that serves as a learning objective for $p_{\boldsymbol{\theta}}$. The approximation gap $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})-\max_{\boldsymbol{\phi}}\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ equals precisely $\min_{\boldsymbol{\phi}}\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$, which motivates the design of expressive classes of posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that reduce this gap. Recent efforts leverage modern generative models—including normalizing flows (rezende2015variational; kingma2016improved) and generative adversarial networks (goodfellow2014generative; makhzani2015adversarial)—as expressive model families for $q_{\boldsymbol{\phi}}$ that tighten the ELBO.
This work seeks to further improve variational inference via expressive posteriors based on diffusion models (ho2020denoising; song2020score). Diffusion methods have become the de-facto standard for high-quality image synthesis (rombach2022high; gokaslan2024commoncanvas). Here, we use diffusion in latent space to parameterize $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$. We train this distribution with a denoising diffusion-like objective that does not involve adversarial training (makhzani2015adversarial) or constrained invertible normalizing flow architectures (kingma2016improved). Samples from $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ are obtained via iterative refinement of $\mathbf{z}$, starting from a Gaussian distribution, and gradually forming one that is multi-modal and complex.
Our work expands upon existing diffusion-based approximate inference methods (berner2022optimal; zhang2021path; vargas2023denoising; zhang2023diffusion; richter2023improved; sendera2024diffusion; akhound2024iterated) that focus on the task of drawing samples from unnormalized distributions $\tilde{p}(\mathbf{z})$ and estimating the partition function $Z=\int_{\mathbf{z}}\tilde{p}(\mathbf{z})d\mathbf{z}$. While these methods are applicable in our setting—we set the unnormalized $\tilde{p}(\mathbf{z})$ to $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ such that $Z=p_{\boldsymbol{\theta}}(\mathbf{x})$—they do not make use of characteristics of $p_{\boldsymbol{\theta}}$ that are common in many types of models (VAEs, Bayes networks, etc.), namely the factorization $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and efficient sampling. We find that leveraging these properties yields simpler algorithms that avoid backpropagating through a sampling process, and that are fast enough to perform learning in addition to inference.
Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion and a learning objective inspired by the wake-sleep algorithm (hinton1995wake) that implements regularized variational inference. We also derive extensions of our method to semi-supervised learning and clustering.
Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks.
We evaluate DDVI on synthetic benchmarks and on a real problem in biological data analysis—inferring human ancestry from genetic data. Our method outperforms strong baselines on 1000 Genomes dataset (siva20081000) and learns a low-dimensional latent space that preserves biologically meaningful structure (haghverdi2015diffusion).
Contributions.
In summary, this work introduces denoising diffusion variational inference, an approximate inference algorithm defined by two components: a class of approximate posteriors $q(\mathbf{z}|\mathbf{x})$ parameterized by diffusion, and a lower bound on the marginal likelihood inspired by wake-sleep. We complement DDVI with extensions to semi-supervised learning and clustering. Our method is especially suited for probabilistic programming, representation learning, and dimensionality reduction, where it outperforms alternative methods based on normalizing flows and adversarial training.
Figure 1: Denoising diffusion variational inference in a VAE. Between the encoder and decoder, we have a diffusion model to map a simple distribution into a complex distribution over latents.
2 Background
Deep Latent Variable Models
Latent variable models (LVMs) $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ are usually fit by optimizing the evidence lower bound (ELBO)
$\log p_{\boldsymbol{\theta}}(\mathbf{x})\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z})),$
which serves as a tractable surrogate for the marginal log-likelihood (MLL). The gap between the MLL and the ELBO equals precisely $\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO (kingma2013auto).
Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods (maaloe2016auxiliary), MCMC-based methods (salimans2015mcmc), normalizing flows (rezende2015variational)—or improved learning objectives—e.g., adversarial or sample-based losses (makhzani2015adversarial; zhao2017infovae; si2022autoregressive; pmlr-v202-si23a).
The wake-sleep algorithm (hinton1995wake) optimizes an alternative objective
$\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]-\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})),$
in which the KL divergence term is reversed. The learning procedure for wake-sleep involves alternating between ”wake” phases where the recognition model is updated and ”sleep” phases where the generative model is refined.
Denoising Diffusion Models
A diffusion model is defined via a user-specified noising process $q$ that maps data $\mathbf{x}_{0}$ into a sequence of $T$ variables $\mathbf{y}_{1:T}=\mathbf{y}_{1},...,\mathbf{y}_{T}$ that represent increasing levels of corruption to $\mathbf{x}_{0}$. We obtain $\mathbf{y}_{1:T}$ by applying a Markov chain $q(\mathbf{y}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where we define $\mathbf{y}_{0}=\mathbf{x}_{0}$ for convenience. When $\mathbf{x}_{0}$ is a continuous vector, a standard choice of transition kernel is $q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{I})$, which is a Gaussian centered around a copy of $\mathbf{y}_{t-1}$ to which we added noise following a schedule $0<{\boldsymbol{\alpha}}_{1}<{\boldsymbol{\alpha}}_{2}<...<{\boldsymbol{\alpha}}_{T}=1$.
A diffusion model can then be represented as a latent variable distribution $p(\mathbf{x}_{0},\mathbf{y}_{1:T})$ that factorizes as $p(\mathbf{x}_{0},\mathbf{y}_{1:T})=p(\mathbf{y}_{T})\prod_{t=0}^{T-1}p_{\boldsymbol{\theta}}(\mathbf{y}_{t}\mid\mathbf{y}_{t+1})$ (again using $\mathbf{y}_{0}$ as shorthand for $\mathbf{x}_{0}$). This model seeks to approximate the reverse of the forward diffusion $q$ and map noise $\mathbf{y}_{T}$ into data $\mathbf{x}_{0}$.
The true reverse of the process $q$ cannot be expressed in closed form; as such, we parameterize $p_{\boldsymbol{\theta}}$ with $\boldsymbol{\theta}$} trained by maximizing the ELBO:
$\log p_{\boldsymbol{\theta}}(\mathbf{x}_{0})\geq$ $\mathbb{E}_{q}\Bigr[\log p_{\boldsymbol{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})-\sum_{t=2}^{T}\text{KL}(q_{t}||p_{t})\Bigr]$ (2)
$-\text{KL}(q(\mathbf{x}_{T}|\mathbf{x}_{0})||p(\mathbf{x}_{T}))$
where $q_{t},p_{t}$ denote the distributions $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ and $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, respectively.
3 Variational Inference With
Denoising Diffusion Models
We introduce denoising diffusion variational inference (DDVI), which improves variational inference with diffusion-based techniques.
The goal of DDVI is to fit a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$. We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) (kingma2013auto).
Our approach is comprised of three components:
  1. 1.
    A modeling family of approximate posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ based on diffusion;
  2. 2.
    A learning objective formed by a regularized ELBO;
  3. 3.
    An optimization algorithm inspired by wake-sleep.
The $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ iteratively refines latents $\mathbf{z}$, starting from a Gaussian distribution. The learning objective trains $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to reverse a used-specified forward diffusion process.
3.1 Modeling Family: Diffusion-Based Posteriors
DDVI performs variational inference using a family of approximate posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})=\int_{\mathbf{y}}q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{x})q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})d\mathbf{y}$, which themselves contain latent variables $\mathbf{y}\in\mathcal{Y}$. The models $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{x}),q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})$ must have tractable densities and support gradient-based optimization over $\boldsymbol{\phi}$}.
We choose the latent $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ to be a vector of $T$ variables that represent progressively simplified versions of $\mathbf{z}$, with $\mathbf{y}_{T}$ corresponding to a simple distribution (e.g., a Gaussian). The model $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\boldsymbol{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$ transforms $\mathbf{y}_{T}$ into $\mathbf{z}$ via iterative refinement. To sample from $q_{\boldsymbol{\phi}}$, we first sample $\mathbf{y}_{T}$—this is an easier task since we can define $\mathbf{y}_{T}$ to have a simple (e.g., Gaussian) distribution—and then by sampling from the denoising model $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\boldsymbol{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$.
We define the relationship between $\mathbf{y}$ and $\mathbf{z}$ via a forward diffusion process $r(\mathbf{y}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1}|\mathbf{z},\mathbf{x})\prod_{t=1}^{T-1}r(\mathbf{y}_{t+1}|\mathbf{y}_{t},\mathbf{x})$, which transforms $\mathbf{z}$—the latent whose intractable posterior we seek to approximate—into $\mathbf{y}_{T}$, whose posterior is easier to model (possibly conditioned on $\mathbf{x}$). Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes (austin2021structured). The model $q_{\boldsymbol{\phi}}$ is trained to approximately reverse this forward diffusion process.
3.2 Learning Objective: A Markovian ELBO
The standard approach to fit auxiliary-variable generative models (maaloe2016auxiliary) is to apply the ELBO twice:
$\log p_{\boldsymbol{\theta}}(\mathbf{x})$ $\geq\log p_{\boldsymbol{\theta}}(\mathbf{x})-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$ (3)
$\geq\log p_{\boldsymbol{\theta}}(\mathbf{x})-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$ (4)
$~~~~-\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$
$=~\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]$ (5)
$~~~~-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$
In Equation (3), we applied the ELBO over $\mathbf{z}$, and in Equation (4) we applied the ELBO over the latent $\mathbf{y}$ of $q$ (see Appendix J for the derivation). Notice that the gap between the ELBO and $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ is $\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))+\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$. Thus, if we correctly match $q$ and $r$, we will achieve a tight bound.
Analyzing the ELBO
Optimizing Equation (5) requires tractably dealing with the prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}}):=-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$, which we equivalently rewrite as:
$\mathcal{L}_{\text{reg}}=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log(r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))]+H(q).$ (6)
We can expand the first term by leveraging the Markov structure of $r,q$ to rewrite $\mathcal{L}_{\text{reg}}$ as the likelihood of samples from the reverse diffusion process $q$ under the forward process $r$.
$\mathcal{L}_{\text{reg}}=\sum_{t=1}^{T}\mathbb{E}_{q}[\log(r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x})]+\mathbb{E}_{q}[\log p(\mathbf{z})]+H(q),$
where $\mathbf{y}_{0}:=\mathbf{z}$. We refer to optimizing the Markovian ELBO as unregularized DDVI, the first instance of our method. The noise process $r$ defines prior regularization terms for each $\mathbf{y}_{t}$. This provides extra supervision for learning $q$ in the form of trajectories from latents $\mathbf{y}_{T}$ to $\mathbf{y}_{1}$; this extra supervision helps $q$ fit complex non-Gaussian posteriors.
The term $H(q)=-\sum_{t=1}^{T+1}\mathbb{E}_{q}[\log q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]$ denotes the entropy. For example, when each term $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ is Gaussian, it is computed as
$H(q)=$ $\sum_{t=1}^{T+1}\mathbb{E}_{q}\left[\frac{d}{2}\left(1+\log(2\pi)\right)+\frac{1}{2}\log\left|\Sigma_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x})\right|\right]$
where $d$ is the dimension of $\mathbf{y}$ and we use the notation $\mathbf{y}_{T+1}=\mathbf{x}$. It is also common to leave the variance $\Sigma_{\boldsymbol{\phi}}$ fixed, in which case $H(q)$ is a constant.
3.3 Refining the Objective: A Regularized ELBO
Notice that optimizing $\mathcal{L}_{\text{reg}}$ involves sampling from the approximate reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ to match the true reverse process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$: this is the opposite of diffusion training, where we would sample from $r$ to fit $q$. This type of on-policy learning of $q$ has been studied in the context of approximate inference (zhang2021path); however, it requires backpropagating through $T$ samples, which may hamper training, and it optimizes a mode-covering divergence that may struggle to fit complex $p(\mathbf{z})$.
Adding Wake-Sleep Regularization to the ELBO
We propose to further improve the ELBO via off-policy diffusion-like training. Our new objective is the ELBO in Equation (5) augmented with a regularizer $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$.
$\log p_{\boldsymbol{\theta}}(\mathbf{x})\geq$ $\underbrace{\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / recons. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}$ (7)
$\underbrace{-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}$
$\underbrace{-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))]}_{\text{sleep term $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$}}$
The optimization of the regularizer $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ is similar to the sleep phase of wake-sleep, and closely resembles diffusion model training (see below). As in wake-sleep, $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ is optimized over $\boldsymbol{\phi}$} only, the $\mathbf{x}$ are sampled from $p$.
From Wake-Sleep to Diffusion Regularization
Computing $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ still involves intractable distributions $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}),q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$. To optimize $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$, we introduce another lower bound $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, which we call the denoising diffusion loss (for reasons that will become apparent shortly):
$\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ $=-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))]$ (8)
$=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}[\log{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\boldsymbol{\theta}})$
$\geq\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}\left[\mathbb{E}_{r}[\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{z},\mathbf{x})}]\right]+\bar{H}(p_{\boldsymbol{\theta}})$
$=\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$
In Equation (8), we applied the ELBO with $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ playing the role of the variational posterior over the latent $\mathbf{y}$ in $q_{\boldsymbol{\phi}}$; $\bar{H}(p_{\boldsymbol{\theta}})$ is the expected conditional entropy of $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on $\boldsymbol{\phi}$}.
We can further simplify $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ by leveraging the Markov structure of the forward and reverse processes $r,q$. Recall that each $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ can be a vector of $T$ latents, which we also denote as $\mathbf{y}_{1:T}$, and that $r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x})$, where $\mathbf{y}_{0}=\mathbf{z}$ and also $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}).$
We may use the Markov structure in $q,r$ to rewrite $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ as a sum of $T$ terms, one per Markov step. The derivation is identical to that used to obtain the ELBO of a diffusion model (sohl2015deep), and yields an expression of the same form:
$\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})=$ $~\mathbb{E}_{r}\left[\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}\text{KL}(r_{t}||q_{t})\right]$ (9)
$-\text{KL}(r(\mathbf{y}_{T}|\mathbf{z},\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x}))+\bar{H}(p_{\boldsymbol{\theta}}).$
where $r_{t},q_{t}$ denote the distributions $r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0},\mathbf{x})$ and $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ (see Appendix K for the derivation).
Regularized DDVI Objective
We define the full DDVI objective $\mathcal{L}_{\text{ddvi}}$ to be the sum of the aforementioned terms:
$\mathcal{L}_{\text{ddvi}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})=\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$
Terms $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{diff}}$ may be weighted by hyper-parameters $\beta_{\text{reg}},\beta_{\text{diff}}>0$, as in the $\beta$-VAE framework. In our experiments, $\beta_{\text{reg}}=\beta_{\text{diff}}=1$ unless otherwise specified. Note that since $\mathcal{L}_{\text{diff}}\leq\mathcal{L}_{\text{sleep}}\leq 0$, $\mathcal{L}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ is a valid lower bound on $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ that is tight when $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})=p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$.
3.4 Optimization: Extending Wake-Sleep
We may optimize $\mathcal{L}_{\text{ddvi}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ using gradient descent by alternating between ELBO optimization and taking sleep steps (see Appendix A for full details):
  1. 1.
    Sample $\mathbf{x}$ from data, take gradient step on ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ optimizing $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ (the “wake” step);
  2. 2.
    Sample $\mathbf{z},\mathbf{x}$ from $p_{\boldsymbol{\theta}}$ and take a gradient step on $\boldsymbol{\phi}$} optimizing $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ (the “sleep” step).
Again, terms may be weighted by $\beta_{\text{reg}},\beta_{\text{diff}}>0$. Note that by maximizing $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, we fit $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ via the forward KL divergence; similarly, by optimizing $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{reg}}$ (the ELBO), we fit $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ via the reverse KL divergence. Thus, optimizing $\mathcal{L}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ encourages $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to approximate $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$, and when the two are equal, the bound $\mathcal{L}_{\text{ddvi}}$ on $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ is tight.
Simplified Wake-Sleep
We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z})$ do not depend on $\mathbf{x}$. This scenario admits the following optimization procedure:
  1. 1.
    Sample $\mathbf{x}$ from data and compute gradient on ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ optimizing $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$.
  2. 2.
    Sample $\mathbf{z}$ from $p(\mathbf{z})$ and compute gradient on $\boldsymbol{\phi}$} optimizing $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$; take step on weighted sum of both gradients.
In this case, $\mathcal{L}_{\text{diff}}$ requires only sampling from $p(\mathbf{z})$, and the entire loss $\mathcal{L}_{\text{ddvi}}$ can be optimized end-to-end using gradient descent. This algorithm is simpler (there is no separate sleep phase); however, $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ may not perfectly approximate $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ when $r(\mathbf{y}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$, hence $\mathcal{L}$ may no longer be a tight bound.
Practical Considerations
[…]
4 Extensions
4.1 Semi-Supervised Learning
Following makhzani2015adversarial, we extend our algorithm to the semi-supervised learning setting where some data points have labels denoted by $l$. We assume the user provides a model of the form $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\boldsymbol{\theta}}(\mathbf{z}|l)p(l)$; we set the variational distributions to $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x}),q_{\boldsymbol{\phi}}(l|\mathbf{x})$. In this setting, we consider two cases, depending on whether the label is observed (kingma2014semi). We extend Equation (7) to incorporate the label $l$ corresponding to a data point as follows:
$\mathcal{L}_{\text{semi}}=$ $~\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},l)]$ (10)
$-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)||p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z}|l))$
$-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}\left[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x},l)||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x},l))\right]$
When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\boldsymbol{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{\phi}})+\text{KL}(q_{\boldsymbol{\phi}}(l|\mathbf{x})||p(l))$. Therefore, we can conclude a marginal likelihood on our dataset as follows: $\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
We also want to guarantee that all model parameters can be learned in all cases, including $q_{\boldsymbol{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference. Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: $\tilde{\mathcal{L}}_{\text{semi}_{\boldsymbol{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\boldsymbol{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\boldsymbol{\phi}}(l|\mathbf{x})\right]$
4.2 Clustering
We have further extended our algorithm to encompass the clustering paradigm. We propose two distinct strategies. In the first approach, we simply formulate a model in which $p_{\boldsymbol{{\boldsymbol{\theta}}}}(\mathbf{z})$ is a mixture of desired priors. The means of these priors are characterized by $\boldsymbol{{\boldsymbol{\theta}}}$}. From these means, cluster membership, denoted as $\mathbf{c}$ can be deduced. This approach requires no alteration to the existing learning objective.
Alternatively, the second method retains the original prior, but introduces an additional latent cluster variable $\mathbf{c}$ where $\sum_{i}c_{i}=1$. Thus, the model can be specified as $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{c})=p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})r(\mathbf{y}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})p(\mathbf{c})$ with $p(\mathbf{c})=Dir({\boldsymbol{{\boldsymbol{\epsilon}}}})$. Consequently, the variational distributions become $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{c},\mathbf{x}),q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{c}|\mathbf{x})$. This yields the objective:
$\mathcal{L}_{\text{clus}}(\mathbf{x})=$ $~\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})]$ (11)
$-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z},\mathbf{c}))$
$-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}\left[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))\right]$
Expectations over small numbers of classes $\mathbf{c}$ are done analytically; larger $\mathbf{c}$ require backpropagating through discrete sampling (jang2016categorical; sahoo2023backpropagation).
Figure 2: Unsupervised visualization on MNIST using three priors (pinwheel, swiss roll, and square). Each color indicates a class.
Method Pinwheel Swiss Roll Square ELBO MMD Latent NLL ELBO MMD Latent NLL ELBO MMD Latent NLL AEVB $12.13\pm 0.41$ $0.77\pm 0.04$ $1.68\pm 0.31$ $14.80\pm 0.23$ $0.78\pm 0.17$ $5.65\pm 1.58$ $7.85\pm 0.29$ $1.10\pm 0.66$ $2.78\pm 0.61$ AEVB-IAF $4.19\pm 0.05$ $0.77\pm 0.00$} $1.64\pm 0.73$ $5.10\pm 0.30$ $0.61\pm 0.15$ $4.43\pm 1.09$ $3.97\pm 0.22$ $0.75\pm 0.12$ $1.68\pm 0.27$ AAEB $N/A$ $0.68\pm 0.02$ $1.54\pm 0.19$ $N/A$ $\mathbf{0.52\pm 0.03}$ $3.34\pm 0.16$} $N/A$ $0.80\pm 0.02$ $2.46\pm 0.46$ H-AEVB $7.03\pm 3.13$ $0.74\pm 0.02$} $2.25\pm 3.02$ $7.21\pm 4.62$ $0.70\pm 0.22$} $4.04\pm 4.62$ $5.71\pm 3.05$ $0.76\pm 0.21$} $2.22\pm 2.03$} PIS $7.83\pm 0.64$ $0.75\pm 0.14$ $6.50\pm 1.11$ $9.83\pm 0.61$ $0.61\pm 0.03$ $\mathbf{2.40\pm 1.01}$ $6\pm 0.06 0.77±0$$7\pm 0.04$ $3.67\pm 0.08$ DDVI $\mathbf{-3.88\pm 0.96}$ $\mathbf{0.67\pm 0.04}$ $\mathbf{1.27\pm 0.21}$ $\mathbf{-5.03\pm 0.58}$ $0.62\pm 0.33$ $3.86\pm 0.17$ $\mathbf{-3.79\pm 0.14}$ $\mathbf{0.66\pm 0.07}$ $\mathbf{1.56\pm 0.09}$
Table 1: Unsupervised learning on MNIST. We report ELBO, MMD between generated images and test images, and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
5 Experiments
We compare DDVI with Auto-Encoding Variational Bayes (AEVB) (kingma2013auto), AEVB with inverse autoregressive flow posteriors (AEVB-IAF) (kingma2016improved), Adversarial Auto-Encoding Bayes (AAEB) (makhzani2015adversarial), and Path Integral Sampler (PIS) (zhang2021path) on MNIST (lecun1998mnist) and CIFAR-10 (krizhevsky2009cifar) in unsupervised and semi-supervised learning settings, and also on the Thousand Genomes dataset (siva20081000). We also compare with Hierachical Auto-Encoding Variational Bayes (H-AEVB) (ranganath2016hierarchical; vahdat2020nvae) in unsupervised setting. We discuss the computational costs of all methods in Appendix D. The priors, model architecture, and training details can also be founded in Appendix H. All results below are reported with 95% confidence interval using 3 different seeds.
5.1 Unsupervised learning
We start with synthetic experiments that are aimed at benchmarking the expressivity of diffusion-based posteriors and their ability to improve fitting $p$, a distribution with a complex structured prior, like one might find in probabilistic programming, scientific analysis, or other applications. We fit a model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ on the MNIST and CIFAR-10 datasets with three priors $p(\mathbf{z})$: pinwheel, swiss roll, and square and report our results in Table 1 and 7. The model distribution $p_{\boldsymbol{\theta}}$ is instantiated by a deep Gaussian latent variable model (DGLVM) with multi-layer perceptrons (MLPs) on MNIST and convolutional neural networks (CNNs) on CIFAR-10 (see the details of model architecture in Appendix G).
Our first set of metrics (ELBO and MMD) seeks to evaluate the learned generative model $p_{\boldsymbol{\theta}}$ is good. In the ELBO calculation, we average the reconstruction loss across image pixels. We use MMD to measure sample quality: we generate images with the trained model and calculate MMD between the generated images and test images using a mixture of Gaussian kernel with sigma equal to $[2,5,10,20,40,80]$. We only report MMD for MNIST, since CIFAR-10 generated samples are very low-quality for all methods because the latent dimension is 2.
Our last metric seeks to directly evaluate the expressivity of the posterior. We measure latent negative log-likelihood (Latent NLL) by fitting a kernel density estimator (KDE) on the latents produced by the model with test data as input and compute the log-likelihood of the latents sampled from the prior under the fitted KDE.
From Table 1 and Table 7 in Appendix, we see our method DDVI achieve best ELBO in all but one scenario, in which it still performs competitively. We also see strong results in Latent NLL and k-nearest neighbors classification accuracy of the latents (Acc). in many scenarios, except for swiss roll where AAEB does well. We present visualizations on MNIST using the baselines and our method in Figure 2.
Figure 3: Semi-supervised visualization on MNIST with 1,000 labels using three different priors (pinwheel, swiss roll, and square). Each a indicates one class.
Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $11.15\pm 0.53$ $0.93\pm 0.01$ $1.36\pm 0.03$ $15.29\pm 1.33$ $0.68\pm 0.01$ $4.60\pm 0.23$ $10.26\pm 0.25$ $0.86\pm 0.01$ $1.68\pm 0.02$ AEVB-IAF $2.10\pm 0.26$ $\mathbf{0.95\pm 0.00}$ $\mathbf{1.06\pm 0.03}$ $5.38\pm 1.78$ $0.90\pm 0.02$ $2.75\pm 0.14$ $2.67\pm 0.83$ $\mathbf{0.91\pm 0.01}$ $\mathbf{0.90\pm 0.02}$ AAEB $N/A$ $0.89\pm 0.01$ $1.55\pm 0.01$ $N/A$ $0.88\pm 0.01$ $3.07\pm 0.05$ $N/A$ $1.94\pm 0.38$ $0.76\pm 0.13$ DDVI $\mathbf{-0.24\pm 0.13}$ $\mathbf{0.95\pm 0.00}$ $\mathbf{1.06\pm 0.01}$ $\mathbf{-2.89\pm 0.33}$ $\mathbf{0.92\pm 0.01}$ $\mathbf{2.09\pm 0.00}$ $\mathbf{0.02\pm 0.09}$ $0.90\pm 0.01$ $1.49\pm 0.03$
Table 2: Semi-supervised learning on MNIST (1,000 labels). We report ELBO, accuracy using KNN (K=20) classifier (Acc), and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
5.2 Semi-supervised Learning
We also evaluate the performance of our method and the baselines under semi-supervised learning setting where some labels are observed (1,000 for MNIST and 10,000 for CIFAR-10) and the partitions of the priors are known.
For this setting, we evaluate ELBO, Latent NLL, and Acc. We choose classification accuracy since classification is a common downstream task for semi-supervised learning. We use the same set of priors and baselines. Details on how we partition each prior into $p_{\boldsymbol{{\boldsymbol{\theta}}}}(\mathbf{z}|\mathbf{x},l)$ can be founded in Appendix F. The partitions defined for our priors are local parts of the priors. We note that unlike unsupervised learning, we use the simplified sleep term in our objective for this setting (see Appendix B for details), since $q_{\boldsymbol{\phi}}$ already gets extra information from $l$ here.
The results are shown in Table 2 and Table 8 in Appendix. DDVI mostly outperforms the baselines across different priors and metrics, especially on CIFAR-10 where DDVI is best across the board. For MNIST, DDVI always achieves the best ELBO, and it also performs competitively with other baselines in classification accuracy. We also show the visualizations of the latents in Figure 3 where DDVI matches the prior almost perfectly.
Method Cluster Purity Cluster Completeness NMI
AEVB $0.28\pm 0.02$ $\mathbf{0.78\pm 0.16}$ $0.59\pm 0.08$
AEVB-IAF $0.29\pm 0.04$ $\mathbf{0.73\pm 0.06}$ $0.55\pm 0.06$
AAEB $0.37\pm 0.06$ $\mathbf{0.76\pm 0.11}$ $0.63\pm 0.02$}
DDVI $\mathbf{0.45\pm 0.03}$ $\mathbf{0.75\pm 0.05}$ $\mathbf{0.66\pm 0.04}$
Table 3: Quantitative genotype clustering results.
5.3 Clustering and Visualization for Genotype Analysis
Figure 4: Visualization of genotype clusters. A color represents one ethnicity.
In this section, we report results on an real-world task in genome analysis. Visualizing genotype data reveals patterns in the latent ancestry of individuals. We instantiate DDVI with a deep Gaussian latent variable model (DGLVM) and compare it against with the three strong clustering baselines using the 1000 Genomes dataset. We also report visualizations from three dimensionality reduction algorithms: PCA, TSNE, and UMAP. For each clustering algorithm, we seek to discover up to 20 clusters. We report quantitative results in terms of cluster purity, cluster completeness, and normalized mutual information (NMI). There is an inherent trade-off between cluster purity completeness. The overall clustering performance can be captured with NMI.
In Table 3, we see that DDVI attains the best performance on cluster purity and NMI. For cluster completeness, VAE and AAE have better means but much larger confidence interval. Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA (wold1987principal), t-SNE (van2008visualizing), and UMAP (mcinnes2018umap)). The legend of Figure 4 can be founded at Figure 5 in Appendix.
6 Discussion
Diffusion vs. Normalizing Flows
Our approach is most similar to flow-based approximators (rezende2015variational; kingma2016improved); in fact when $T\to\infty$, our diffusion-based posterior effectively becomes a continuous-time normalizing flow (song2020denoising). However, classical flow-based methods require invertible architectures for each flow layer: this constrains their expressivity and requires backpropagating through potentially a very deep network. Our approach, on the other hand, trains a model (a continuous-time flow when $T\to\infty$) via a denoising objective (similar to score matching) that does not require invertible architectures and effectively admits an infinite number of layers (with weight sharing). This model is trained not by backpropagating through the ELBO, but rather via an auxiliary diffusion loss term (effectively, a score matching objective).
Despite training with a modified loss, we observe in Section 5 that a diffusion model with an expressive denoising architecture yields an improved ELBO relative to regular flows. Also, our modified loss based on the forward KL divergence reduces posterior collapse (i.e., all modes of the prior are covered well), and thus produces better samples.
Diffusion vs. Other Generative Models
Variational posteriors based on GANs (makhzani2015adversarial) also admit expressive architectures and require only sample-based access to the prior $p(\mathbf{z})$. Our diffusion-based approach admits a more stable loss, and is potentially more expressive, as it effectively supports an infnite number of layers (with shared parameters when $T\to\infty$). Unlike GANs, our models also admit explicit likelihoods and allow us to compute the ELBO for model evaluation. Our approach is similar to variational MCMC (salimans2015mcmc); however, we train with a better objective augmented with a diffusion loss, and we adopt improved architectures with shared weights across layers.
Diffusion for Approximate Inference
Existing diffusion-based approximate inference methods (berner2022optimal; zhang2021path; vargas2023denoising; zhang2023diffusion; richter2023improved; sendera2024diffusion; akhound2024iterated) focus on the task of drawing samples from unnormalized distributions $\tilde{p}(\mathbf{z})$ and estimating the partition function $Z=\int_{\mathbf{z}}\tilde{p}(\mathbf{z})d\mathbf{z}$. While these methods are applicable in our setting—we set the unnormalized $\tilde{p}(\mathbf{z})$ to $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ such that $Z=p_{\boldsymbol{\theta}}(\mathbf{x})$—they also tackle a more challenging problem (drawing samples from energy-based models) in more general classes of models (arbitrary unnormalized distributions). In contrast, we focus on restricted but still important classes of models (VAEs, Bayes networks, etc.), and we solve more challenging sets of tasks (e.g., maximum-likelihood learning) by using properties of $p_{\boldsymbol{\theta}}$ (the factorization $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and efficient sampling from $p_{\boldsymbol{\theta}}$).
Our algorithms are also simpler. For example, diffusion sampling methods require backpropagating through a sampling process to minimize the reverse $\text{KL}(q_{\boldsymbol{\phi}}||p_{\boldsymbol{\theta}})$, which poses challenges with optimization and credit assignment. Some methods based on Schrodinger bridges require an iterative optimization process generalizing the sinkhorn algorithm or computationally expensive on-policy or off-policy (malkin2022trajectory) trajectory-based optimization. In contrast, DDVI optimizes the forward $\text{KL}(p_{\boldsymbol{\theta}}||q_{\boldsymbol{\phi}})$ using simple gradient-based optimization that directly emulates diffusion-based training.
7 Related Work
Latent Diffusion
vahdat2021score; wehenkel2021diffusion; rombach2022high perform diffusion in the latent space of a VAE. Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors.
Our goal is different: we seek to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$. Recent work (preechakul2022diffusion; zhang2022unsupervised; wang2023infodiffusion) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$. cohen2022diffusion crafts a diffusion bridge linking a continuously coded vector to a non-informative prior distribution.
Diffusion for Approximate Inference
Diffusion sampling (berner2022optimal; zhang2021path; vargas2023denoising; zhang2023diffusion; richter2023improved; sendera2024diffusion; akhound2024iterated) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function. These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift). Some other works (zhang2023diffusion; akhound2024iterated; sendera2024diffusion) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration. While this work is applicable to our setting, it does not rely on the structure of $p(\mathbf{x},\mathbf{z})$ available to us, namely tractable sampling in $p$.
Dimensionality Reduction
Latent variable models in general are an attractive alternative to visualization methods like PCA, Sparse PCA, NMF, UMAP, and t-SNE (wold1987principal; kuleshov2013fast; NIPS2000_f9d11525; mcinnes2018umap; van2008visualizing). Domain-specific knowledge can be injected through the prior, and deep neural networks can be utilized to achieve a more expressive mapping from the data space to the latent space. Nevertheless, downsides of LVMs are that they are more computationally expensive and require careful hyperparameter tuning.
8 Conclusion
While this paper focuses on applications of DDVI to dimensionality reduction and visualization, there exist other tasks for the algorithm, e.g., density estimation or sample quality. Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making (DBLP:conf/icml/NguyenG22; deshpande2023calibrated), meta-learning (rastogi2023semi), or causal effect estimation (deshpande2022deep).
Since our learning objective differs from the ELBO (it adds a regularizer), we anticipate gains on models whose training benefits from regularization, but perhaps not on all models. Also, attaining competitive likelihood estimation requires architecture improvements that are orthogonal to this paper. However, our ability to generate diverse samples and achieve class separation in latent space hints at the method’s potential on these tasks.
Appendix
Appendix A Pseudocode
Here we provide a pseudocode to illustrate the training process of DDVI.
Algorithm 1 DDVI Pseudocode
1:(Optional) Pre-train $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})$ with DDVI but with unconditional diffusion model $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$
2:for epoch $1,\ldots,n$ do
3:  for $\mathbf{x}_{1},\ldots,\mathbf{x}_{k}\sim p_{\mathcal{D}}(\mathbf{x})$ do
4:   $\mathbf{y}_{i}\sim q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x}_{i})$ and $\mathbf{z}_{i}\sim q(\mathbf{z}|\mathbf{y}_{i},\mathbf{x}_{i})$ for $i=1,\ldots,k$
5:   Optimize ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ with respect to a Monte Carlo estimate of $\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\|p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z})\right)$ for each $\mathbf{x}_{i}$ $\triangleright$ Standard ELBO training part
6:   for iteration $1,\ldots,m$ do $\triangleright$ Do sleep for $m$ iterations
7:     $\mathbf{z}_{1},\ldots,\mathbf{z}_{k}\sim p(\mathbf{z})$ $\triangleright$ Batch-sample latents from prior
8:     $\hat{\mathbf{x}}_{i}\sim p(\mathbf{x}|\mathbf{z}_{i})$ for $i=1,\ldots,k$ $\triangleright$ Construct fantasy inputs
9:     $\mathbf{y}_{i}\sim r(\mathbf{y}|\mathbf{z}_{i})$ for $i=1,\ldots,k$ $\triangleright$ Construct fantasy inputs
10:     Optimize $\boldsymbol{\phi}$} using the standard diffusion noise prediction loss on $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{i},\hat{\mathbf{x}}_{i})$
11:   end for
12:  end for
13:end for
Appendix B Simplifying Wake-Sleep
In wake-sleep, sampling $\mathbf{x}$ from $p_{\boldsymbol{\theta}}$ to obtain gradients for the sleep term introduces computational overhead. To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation $\hat{\mathcal{L}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ of $\mathcal{L}$:
[…]
We have replaced $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ with a latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$, in which $\mathbf{x}$ is given, and we only seek to fit the true reverse noising process $r(\mathbf{z}|\mathbf{y})$ independently of $\mathbf{x}$. We can similarly show that
$\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$ $=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{z})}[\log{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\boldsymbol{\theta}})\geq\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{z})r(\mathbf{y}|\mathbf{z})}[\log({q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}/r(\mathbf{y}|\mathbf{z}))]+\bar{H}(p_{\boldsymbol{\theta}})$ (13)
$=-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{z})}[\text{KL}(r(\mathbf{y}|\mathbf{z})||q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{z},\mathbf{x}))]-\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z})||q(\mathbf{z}|\mathbf{x})),$ (14)
where $\bar{H}(p_{\boldsymbol{\theta}})$ is an entropy term constant in $\boldsymbol{\phi}$}. Thus, we minimize the forward KL divergence by sampling $\mathbf{z}$, and applying the noising process to get $\mathbf{y}$; the $q_{\boldsymbol{\phi}}$ is fit to denoise $\mathbf{z}$ from $\mathbf{y}$ as in Equation 8.
We optimize our bound on $\hat{\mathcal{L}}(x,{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ end-to-end using minibatch gradient descent over ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$. While the wake term is a reconstruction loss as in wake-sleep, the sleep term generates latent samples $\mathbf{z},\mathbf{y}$ from $r(\mathbf{y}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ (by analogy with $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ in normal wake-sleep); the denoiser $q_{\boldsymbol{\phi}}$ is trained to recover $\mathbf{z}$ from $\mathbf{y}$. Thus, we perform wake-sleep in latent space, which obviates the need for alternating wake and sleep phases, and allows efficient end-to-end training. A limitation of this approximation is that the sleep term does not fit $q_{\boldsymbol{\phi}}$ to the true $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x},\mathbf{y})$, and as a consequence $\hat{L}$ is not a tight lower bound on $\log p_{\boldsymbol{\theta}}(\mathbf{x})$. We may think of $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$ as a regularizer to the ELBO.
Appendix C Comparision of Methods
We provide a comprehensive comparison of different methods in Table 4. vahdat2021score; wehenkel2021diffusion; rombach2022high perform diffusion in the latent space of a VAE to improve the efficiency of image generation. Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors. Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
Recent work (preechakul2022diffusion; zhang2022unsupervised; wang2023infodiffusion) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$. cohen2022diffusion crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
Model Training Objective Approximating Family Sample-based Prior Auxiliary Variable Tasks Simplified Graphical Illustration AEVB ELBO Diagonal Gaussian Density estimation $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ AEVB-IAF ELBO Normalizing flow Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{x}$ AAEB Adversarial training Adversarial generator Visualization $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ H-AEVB-(IAF) ELBO Factorial Normal / Normalizing flow Density estimation / High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ ADGM ELBO Non-Gaussian Density estimation $\mathbf{x}\rightarrow\mathbf{a}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ LDM ELBO Diagonal Gaussian High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ LSGM ELBO & score matching Diagonal Gaussian High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ DDVI ELBO & sleep term Denoising diffusion Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{T}(\mathbf{y})\rightarrow\mathbf{z}_{0}(\mathbf{z})\rightarrow\mathbf{x}$
Table 4: Comparison of DDVI to other relevant methods. $\mathbf{x}$ represents the original data input to the model. $\mathbf{z}$ denotes the latent (hidden) representation of the input data. $\mathbf{a}$ represents an auxiliary variable introduced in some models (like ADGM) to capture additional aspects of the data distribution or to assist in the model’s learning process.
Appendix D Computational Cost Analysis
Method NMI values at different wall-clock training times NMI @ 10 min NMI @ 20 min NMI @ 30 min NMI @ 40 min NMI @ 50 min NMI @ 60 min AEVB $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ AEVB-IAF $0.54$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ AAEB $0.61$ $0.57$ $0.57$ $0.57$ $0.57$ $0.57$ DDVI (T=5) warm up $0.63$ $0.63$ $0.66$ $0.66$ $0.66$ DDVI (T=10) warm up $\mathbf{0.64}$ $\mathbf{0.68}$ $\mathbf{0.70}$ $\mathbf{0.70}$ $\mathbf{0.70}$ DDVI (T=20) warm up $0.50$ $0.51$ $0.56$ $0.64$ $0.68$ DDVI (T=50) warm up $0.52$ $0.54$ $0.51$ $0.59$ $0.59$
Table 5: Computational cost trade-off on 1kgenome: NMI vs wall-clock training time
We conduct a computational cost analysis between the baselines and DDVI with various timesteps on the genotype clustering/visualization experiments. Table 5 shows that DDVI outperforms baselines at all timestamps and continues to improve after the baselines have plateaued.
Appendix E Connections to Diffusion Samplers
Diffusion sampling (berner2022optimal; zhang2021path; vargas2023denoising; zhang2023diffusion; richter2023improved; sendera2024diffusion; akhound2024iterated) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function. These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) (zhang2021path), denoising diffusion sampler (DDS) (vargas2023denoising), and time-reversed diffusion sampler (DIS) (berner2022optimal), which have been unified by richter2023improved. Some other works (zhang2023diffusion; akhound2024iterated) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration. sendera2024diffusion benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps. Overall, there are indeed some strong connections between these works and ours:
  • They also focus on variational methods that directly fit a parametric family of tractable distributions (given by controlled SDEs) to the target density.
  • They cast the density estimation/sampling problem into an optimization problem over a control objective, which learns control drifts (and diffusion) parameterized by neural networks.
But we would like to clarify that there are also some clear differences between them:
  • The diffusion-structured samplers only focus on density estimation/sampling but ignore the problem of learning a generative model, which is one of the main focuses of our work. We aim to perform more accurate variational inference using an auxiliary variable model augmented by diffusion models to improve generative modeling. In our setting, $p_{\boldsymbol{\theta}}(z|x)$ is a moving target density, as we jointly learn $\boldsymbol{\theta}$} with $\boldsymbol{\phi}$}, as opposed to a static target density that diffusion-structured samplers are designed to solve.
  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy (zhang2021path) or off-policy (malkin2022trajectory) trajectory-based optimization. Instead, we introduce a wake-sleep optimization algorithm and its simplified version to alleviate the weak learning signal issue and optimize the evidence lower bound in a better way.
  • In Equation 8, we are minimizing the forward KL divergence $\text{KL}(p_{\boldsymbol{\theta}}||q_{\boldsymbol{\phi}})$, where diffusion samplers are minimizing the reverse $\text{KL}(q_{\boldsymbol{\phi}}||p_{\boldsymbol{\theta}})$.
We also summarize the connections and differences in the table below.
Diffusion Samplers GFlowNet-based DDVI (ours)
(berner2022optimal; zhang2021path; vargas2023denoising; akhound2024iterated) (zhang2023diffusion; richter2023improved; sendera2024diffusion; malkin2022trajectory)
Tasks Sampling, density estimation Sampling, density estimation Learning, Sampling, Dimensionality reduction
Model Family for $p$ Any energy-based Any energy-based Latent with tractable $p(x|z),p(z)$
Model Family for $q$ Markov chain Markov chain Markov chain
Objective $\text{KL}(q||p)$ with regularizer Trajectory balance objective $\text{KL}(p||q)$ with ELBO
Algorithm Gradient descent (with reference process), importance sampling RL-motivated off-policy optimization (replay buffers, Thompson sampling, etc.) Gradient descent with wake-sleep
Compatible Models Anything energy-based Anything energy-based LDA, deep latent-variable models
Applications Sampling from physics-based models, model selection based on NLL Sampling from physics-based models, model selection based on NLL Probabilistic programming, visualization
Table 6: Comparison of Diffusion-structured Samplers, GFlowNet-based Approaches, and DDVI
Appendix F Priors
Below we describe the sampling process for each prior.
Pinwheel. This distribution was used in (johnson2016svae). We define the number of clusters to be 10. For semi-supervised learning experiments, this prior is partitioned into 10 partitions, each partition being a cluster.
Swiss Roll. This distribution was used in (marsland2014swissroll). For semi-supervised learning experiments, this prior is partitioned into 10 partitions. The samples from the prior can actually be characterized by a single scalar representing how far you are long the swiss roll from the center. The paritioning is done by creating 10 equal-length intervals in this 1D space.
Square. This distribution has the shaped of a square going from -1 to 1 in both axes. Each position on the square can be characterized by a single scalar representing how far you are from the top left corner. Sampling is done by sampling the position uniformly and turn the 1D position to 2D latent. We add noise $\sigma=0.06$ to the prior. For semi-supervised learning experiments, this prior is partitioned into 10 partitions. The partitioning is done by creating 10 equal-length intervals in the 1D position space.
AEVB and AEVB-IAF requires that we can evaluate the prior density. To do this, for all priors, we evaluate the density by fitting a kernel density estimator with mixture of gaussian kernel with bandwidth equal to 0.005, 0.008, 0.01, 0.03, and 0.05.
Appendix G Model Architecture
All methods use the same architecture for encoder $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ and decoder $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$, excluding the extra parts specific to each method which we describe below, for the same dataset. For MNIST, the encoder and decoder are multi-layer perceptron with two hidden layers, each with 1000 hidden units. For CIFAR-10, the encoder is a 4-layer convolutional neural network with (16, 32, 64, 128) channels with a linear layer on top, and the decoder is a 4-layer tranposed convolutional neural network with (64, 32, 16, 3) channels where a linear layer is used to first turn the feature dimension from 2 to 64.
AEVB-IAF employs 4 IAF transformations on top of the encoder, each is implemented with a 4-layer MADE. The number of hidden units in MADE is 128. The ordering is reversed between every other IAF transformation.
AAEB has a discriminator, used in adversarial training, which is a multi-layer perceptron with two hidden layers, each with 1000 hidden units.
DDVI has a diffusion model on top of the encoder. The time-conditioned reverse diffusion distribution is implemented with a 5-layer time-conditioned multi-layer perceptron, each with 128 hidden units. A time-conditioned linear layer learns an additional embedding for each timestep and adds it to the output of the linear layer.
Appendix H Training Details
For training, we update the parameters for each batch of inputs by alternating between the ELBO phase (optimizing $\boldsymbol{\theta}$} and $\boldsymbol{\phi}$} with respect to the ELBO, i.e., the reconstruction term and the prior matching term) and the sleep phase (optimizing $\boldsymbol{\phi}$} with respect to the sleep term). We use Adam optimizer and latent size of 2 for all of our experiments. Each algorithm takes roughly 2 hours on a single Nvidia GeForce RTX 3090 to complete one run of experiment. The training details of each algorithm are detailed below:
AEVB. The batch size is set to 128. The number of epochs is 200 for unsupervised and clustering experiments and 50 for semi-supervised experiments. The learning rate is 0.0001. The loss is BCE for MNIST and CIFAR-10 experiments and MSE for genotype analysis experiments. For semi-supervised MNIST experiments, the kl divergence weight is set to be 0.01, while for semi-supervised CIFAR-10 experiments, the kl divergence weight is set to be 0.01. For other experiments, the KL divergence weight is set with a schedule linear on number of epochs going from 0 to 0.01. We also have a weight of 5 multiplied to the prior density.
AEVB-IAF. The batch size, number of epochs, learning rate, loss, KL divergence weight, and prior density weight are the same as VAE. The context size, i.e., the size of features used to initialize the flow layers for different datat point, is 10.
AAEB. The batch size is set to 128. The number of epochs is 200 for all experiments. The learning rate is 0.0002. The loss is MSE for all experiments. To stabilize the training, we add noise to the input to the discriminator with sigma 0.3 at the start and lower it by 0.1 for every 50 epochs. The noise equals to 0 at epoch 150.
DDVI. The batch size is set to 128 for most experiments, except for semi-supervied experiments where the batch size is 1024. The number of epochs is 200 for unsupervised and clustering experiments and 30 for semi-supervised experiments. The learning rate is 0.0001. The loss is BCE for MNIST and CIFAR-10 experiments and MSE for genotype analysis experiments. For unsupervised MNIST and CIFAR-10 experiments, the KL divergence weight is set to 0.003. For semi-supervised MNIST experiment, we use KL divergence weight of 0.1. For semi-supervised CIFAR-10 experiment, we use KL divergence weight of 0.5. For clustering experiment, we use KL divergence weight of 0.005. The number of timesteps is 20 for unsupervised and clustering experiments and 100 for semi-supervised experiments.
Appendix I Genotype Analysis Experiments Details
Figure 5: Legend showing what ethnicity each color corresponds to in the 1000 Genomes dataset
Before inputting the data points into any of the visualization methods, we first pre-process it by running a PCA and keep only the first 1000 principal components of the data points. We further divide the features by 30 for all latent variables model methods.
The legend of the 1000 Genomes Visualization plot can be found at Figure 5.
Method Pinwheel Swiss Roll Square
ELBO Latent NLL ELBO Latent NLL ELBO Latent NLL
AEVB $12.96\pm 1.81$ $3.26\pm 0.60$ $12.87\pm 4.55$ $6.25\pm 1.58$ $7.91\pm 0.11$ $2.91\pm 0.17$
AEVB-IAF $3.24\pm 0.16$ $1.71\pm 0.84$ $4.03\pm 0.73$ $5.51\pm 0.51$ $\mathbf{-2.10\pm 0.31}$ $1.71\pm 0.77$
AAEB $N/A$ $1.70\pm 0.41$ $N/A$ $\mathbf{3.18\pm 0.22}$ $N/A$ $1.67\pm 0.17$
H-AEVB $4.42\pm 0.46$ $\mathbf{1.69\pm 0.17}$ $5.36\pm 0.77$ $5.74\pm 0.55$ $2.86\pm 0.11$ $1.64\pm 0.09$
PIS $2.92\pm 1.23$} $3.61\pm 0.62$} $4.14\pm 0.49$ $7.14\pm 0.14$ $4.85\pm 0.06$ $3.91\pm 0.06$
DDVI $\mathbf{-1.38\pm 0.44}$ $1.75\pm 0.53$} $\mathbf{-3.05\pm 0.65}$ $5.66\pm 2.63$} $2.47\pm 0.30$ $\mathbf{1.58\pm 0.09}$
Table 7: Unsupervised learning on CIFAR-10. We report ELBO and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\mathbf{-1.60\pm 0.29}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.09\pm 0.05}$ $\mathbf{-4.13\pm 1.51}$ $\mathbf{0.47\pm 0.09}$ $\mathbf{2.29\pm 0.08}$ $\mathbf{-1.73\pm 0.64}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.48\pm 0.02}$
Table 8: Semi-supervised learning on CIFAR-10 (10,000 labels). We report ELBO, accuracy using KNN (K=20) classifier (Acc), and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB - - - H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\mathbf{1.27\pm 0.21}$ $\mathbf{3.86\pm 1.17}$ $\mathbf{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
Table 9: Unsupervised learning on MNIST, including the results of DDVI without the sleep term.
Appendix J ELBO for Auxiliary-Variable Generative Models
We aim to derive the lower bound for the log-likelihood $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ by introducing auxiliary variables and applying the Evidence Lower Bound (ELBO) twice.
Introduce the latent variable $\mathbf{z}$ and apply the ELBO:
$\log p_{\boldsymbol{\theta}}(\mathbf{x})$ $=\log\int p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})\,d\mathbf{z}$ (15)
$\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\right]\quad\text{(by Jensen's inequality)}$ (16)
$=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\text{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\right)$ (17)
Then, introduce an auxiliary variable $\mathbf{y}$ and apply the ELBO again:
$\log p_{\boldsymbol{\theta}}(\mathbf{x})$ $\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\text{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\right)$ (19)
$=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})+\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\log p(\mathbf{z})$ (20)
$\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]\right]-\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})}\left[\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{x},\mathbf{z})}\right]\right]+\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})}\left[\log p(\mathbf{z})\right]\right]$ (21)
$=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\text{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\,\|\,r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z})\right)$ (22)
Appendix K Diffusion Regularization
We begin with the definition of $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, derived as a lower bound on $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$:
$\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}\left[\mathbb{E}_{r}\left[\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{z},\mathbf{x})}\right]\right]+\bar{H}(p_{\boldsymbol{\theta}}).$
To simplify $\mathcal{L}_{\text{diff}}$, we leverage the Markov structure of the forward process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ and the reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$.
The forward process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ is decomposed as:
$r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x}),$
where $\mathbf{y}_{0}=\mathbf{z}$.
The reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ is decomposed as:
$q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}),$
where $\mathbf{y}_{0}=\mathbf{z}$ and $\mathbf{y}_{1:T}$ are increasingly noisy versions of $\mathbf{z}$.
Substituting the factorizations of $r$ and $q_{\boldsymbol{\phi}}$ into the definition of $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, and rewriting the logarithm and rearranging terms, we have:
$\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ $=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}\Bigg[\mathbb{E}_{r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})}\Bigg[\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})}{\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x})}\Bigg]\Bigg]+\bar{H}(p_{\boldsymbol{\theta}})$
$=\mathbb{E}_{r(\mathbf{y}_{1:T},\mathbf{z},\mathbf{x})}\left[\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}\text{KL}(r_{t}||q_{t})\right]-\text{KL}(r(\mathbf{y}_{T}|\mathbf{z},\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x}))+\bar{H}(p_{\boldsymbol{\theta}})$
Here, $\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})$ corresponds to the reconstruction term for the initial latent state $\mathbf{z}$, while the summation represents the KL divergence between the forward and reverse processes at each intermediate step. $r_{t}=r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{z},\mathbf{x})$ is the conditional forward distribution at step $t$, $q_{t}=q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ is the reverse process distribution at step $t$. $\bar{H}(p_{\boldsymbol{\theta}})$ is the expected conditional entropy of $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on $\boldsymbol{\phi}$}.
This form mirrors the ELBO derivation for diffusion models (sohl2015deep), where each step in the Markov chain contributes a KL divergence term, and the reconstruction term arises from the connection between the noisy latent variables and the original data.